# SMMAN: quasi-Simultaneous Multi-wavelength Monitoring of gamma-ray-loud AGNs with the Nanshan 26-m radio telescope

### v1

@@ I Introduction
A few of the major radio band surveys are briefly introduced below, with emphasis on the importance of multi-frequency radio monitoring projects. The University of Michigan Radio Astronomy Observatory’s (UMRAO; Aller et al. 1985) 26-m paraboloid telescope was devoted to monitoring the flux density and polarization of blazar jets from the mid-1960s until June 2012 (Aller et al., 2017). This facility provided centimeter-band variability over decades at 14.5, 8.0, and 4.8 GHz. The UMRAO data have been utilized in numerous extensive studies concerning flaring blazars and to acquire information regarding various jet and shock parameters that are not directly observable (Aller et al., 2014; Hughes et al., 2015; Aller et al., 2016) and also in several multiband temporal and spectral studies (e.g. Krishna Mohana et al., 2024; Mohana A et al., 2025). UMRAO observations were also used for the calibration and interpretation of the less frequently sampled Very Long Baseline Array (VLBA) imaging data (Aller et al. 2017, and references therein). It is to be noted that the long-term monitoring data set could help to search for a supermassive black hole binary (SMBHB)
@@
The observations were carried out with the six frequency bands at 1.25, 2.25, 4.7, 8.2, 11.2, and 22.3 GHz, simultaneously (within 1–2 min) when a source moves along the focal line where the receivers are located. These data formed the basis for the first version of the BLcat (Mingaliev et al., 2014), and the latest updated version of this catalog is “RATAN-600 multi-frequency catalog of blazars” (Sotnikova et al., 2022), which includes about 1830 source. Using these catalogs with information of flux densities, spectral indices, variability, and radio luminosity for the large blazar list, several studies to understand the flare activity and synchrotron radio spectra of AGNs were reported (e.g. Mufakharov et al., 2015; Sotnikova et al., 2019; Larionov et al., 2020; Sotnikova et al., 2021; Krishna Mohana et al., 2024; Kunert-Bajraszewska et al., 2025). The Fermi-GST AGN Multi-frequency Monitoring Alliance (F-GAMMA) programme (Fuhrmann et al., 2007; Angelakis et al., 2010; Fuhrmann et al., 2014), conducted at the Effelsberg telescope and IRAM 30 m telescope in Germany, performed nearly monthly broadband radio monitoring of selected blazars from January 2007 to January 2015 (Angelakis et al., 2019). The monitoring was carried out at multi-frequency from 2.64 to 43 GHz, which includes the frequencies of 2.64, 4.85, 8.35, 10.45, 14.6, 23.05, 32, 43, 86.2, 142.3, and 228.9 GHz. Using initial 3.5 yr data, a cross-correlation analysis was carried out by Fuhrmann et al. (2014) between the radio and -ray light curves of 54 Fermi-bright blazars, with the aim of determining the exact location of the -ray emission region. Fuhrmann et al. (2016) reported a multi-frequency study of AGNs in Fermi era to examine whether the -ray loudness is related to the radio variability. The 2007-2015 multi-frequency radio data published in Angelakis et al. (2019) is used in several temporal and spectral studies of blazars (e.g. Paliya et al., 2021; Kim et al., 2022; Krishna Mohana et al., 2024).
@@
The Radio Observations of Blazars with INaf telescopes (ROBIN) program has continued to thrive over the years and remains active at frequencies of 8 and 24 GHz, having reached 20 years of operation (Marchili et al., 2025). The data accumulated through the program have been utilized in numerous studies and publications, primarily in relation to Whole Earth Blazar Telescope (WEBT) multi-frequency campaigns (e.g. Raiteri et al., 2007; Abdo et al., 2010b; Carnerero et al., 2015; Raiteri et al., 2017, 2024).
The monitoring program has facilitated the examination of how the variability characteristics and SED of blazars evolve over time (Marchili et al. 2025, and references therein).
At the Aalto University Metsähovi Radio Observatory (MRO, Finland), several hundred AGNs and a few galactic microquasars have been monitored at 22, 37, and 87 GHz since 1980 using the 13.7 m radio telescope (Teraesranta et al., 1998). At present, there are 271 sources in the MRO monitoring program with data at 22 and 37 GHz, where the latter includes observations spanning up to 42 years. Data from this ongoing program have contributed to research on the long-term behavior, statistical properties, source populations, and radio variability models of AGNs. This monitoring has enabled studies of the radio to -ray connection in blazars, as well as radio versus optical behavior and multi-frequency variability models (e.g. Hovatta et al., 2007; Kankkunen et al., 2025). The Submillimeter Array (SMA; Gurwell et al., 2007) in Hawaii, United States, has been monitoring the quasars to determine the flux densities and variability at 230 and 345 GHz since the beginning of 2003 with 437 sources at present. The observations from this program were utilized in different long-term timing and spectral studies of AGNs (e.g. Villata et al., 2009; Fuhrmann et al., 2014; Jorstad and Marscher, 2016).
@@ II.1 Facility
The SMMAN program is carried out with the NanShan Radio Telescope (NSRT), which is located on the Nanshan Mountain, with the geographic coordinates of 43°28’17”N, 87°10’41”E, and an altitude of 2080 meters, km southwest of Urumqi city, China. Apart from single-dish scientific observations, NSRT has been involved in a large number of Very Long Baseline Interferometry (VLBI) observations in the last years, as a formal member of the European VLBI Network (EVN). NSRT now consists of a 26-meter parabolic antenna and four receivers operating at L-band (1-2 GHz), S/X-band (2.3/8.6 GHz), C-band (4-8 GHz), and K-band (22-24 GHz). The receivers utilized in the SMMAN program operate in the C and K bands, centered at 4800 MHz and 23600 MHz, with terminal recording bandwidths of 500 MHz for both. The main technical specifications of the receivers used in the SMMAN program are listed in Table 1.
@@ II.2 Source selection and monitoring schedule
Our goal is to monitor those bright variable blazars, the majority of which have been captured by the Fermi-LAT. Therefore, the initial source sample was established by filtering the Fermi-LAT 3FGL catalog (Acero et al., 2015) and cross-matching with the NRAO VLA Sky Survey (NVSS)222https://www.cv.nrao.edu/nvss/ by putting a criteria of flux density at 1.4 GHz (, point-like structure, and source declination (J2000) .
@@ II.3 Data Acquisition and Reduction
The data reduction and calibration process includes Gaussian fitting, pointing correction, gain correction, and absolute flux density calibration. Gaussian fitting is used to subtract the baseline of the system temperature drifting and extract the Gaussian-like signals from the sub-scans. Pointing correction is to calibrate the antenna pointing error with each other direction’s offset in the scans. After that, we subsequently conducted atmospheric opacity and antenna gain calibration to further correct the elevation gain effect. Finally, the absolute flux density was tuned using the calibration factors generated by three primary calibrators, 3C 48, NGC 7027, and 3C 286 (Baars et al., 1977; Ott et al., 1994; Zijlstra et al., 2008; Angelakis et al., 2015). The observed average flux density for these calibrator sources during the SMMAN program is presented in Table 2.
@@
We utilized the monthly binned -ray flux and spectral index data from the Fermi-LAT Light Curve Repository444https://fermi.gsfc.nasa.gov/ssc/data/access/lat/LightCurveRepository/ (LCR; Abdollahi et al. 2023) in the energy range of GeV, covering the same time period as the radio observations (2016 November 1 to 2024 November 30) presented in this study. For a detailed description of the unbinned maximum-likelihood data analysis, we refer to Abdollahi et al. (2020). In this work, we employed the light curve obtained by allowing the spectral slope to vary freely during the fitting process and the fit results of only the convergent likelihood analysis.
The significance of a source detection is evaluated using the test statistic , where is the likelihood of the null model (no source) and is the likelihood of the model that includes the source. In the monthly binned light curve, we only consider those bins with detection significance of (), and the upper limit values (for non-detection bins) were not included. It is important to note that performing a full Fermi-LAT data analysis for 131 SMMAN sources is computationally demanding and time-consuming. Therefore, we used the available light curves from the LCR. These light curves of various -ray sources from this repository have been utilized in previous research studies (e.g. Di Gesu et al., 2023; Wang et al., 2023).
@@ III Results and Discussion
To depict the importance of the SMMAN monitoring program, a long-term multiband light curve of the BL Lac object S4 0954+65 using the data from the SMMAN, RATAN-600555https://www.sao.ru/blcat/(Mingaliev et al., 2014; Sotnikova et al., 2022), F-GAMMA666https://cdsarc.cds.unistra.fr/viz-bin/cat/J/A+A/626/A60 (Angelakis et al., 2019) and ROBIN777https://cdsarc.cds.unistra.fr/viz-bin/cat/J/A+A/702/A140 (Marchili et al., 2025) monitoring program is shown in Figure 5. The plot clearly shows that the SMMAN program has covered a greater number of observations after 2016 November compared to the other surveys (publicly available), which is important due to the fact that adequate continuous data are essential for long-term temporal and spectral studies, and the detection of QPOs. In addition, a part of the data from the SMMAN program has been utilized in systematic studies of a few individual blazars in earlier research works (e.g. Kun et al., 2022; Krishna Mohana et al., 2024; Kun et al., 2024; Mohana A et al., 2025), which also showcase the significance of the SMMAN program along with other multiwavelength observations.
@@ III.2 Characteristics of Radio Variability, Spectral index and Luminosity: Relation to AGNs Class
Table 5 of the Appendix A.
@@ III.3 Gamma-ray and Radio Properties of the SMMAN Sample
A linear fit in log-log space gives (Pearson correlation coefficient = 0.68) and (Pearson correlation coefficient = 0.86) for radio luminosities of 4.8 and 23.6 GHz, respectively. Such a correlation could be induced by observational effects, which depend on the source distance (e.g., Böck et al., 2016).
@@ IV Summary and Conclusions
The users utilizing the SMMAN radio data in their publications are kindly requested to acknowledge the source of the data by citing this paper. To use the data and for details, please contact Lang Cui (cuilang@xao.ac.cn).
@@ Acknowledgements
We are thankful to Sándor Frey, Timur Mufakharov, and Junhui Fan for carefully reading the manuscript and providing insightful comments. This work was supported by the Tianshan Talent Training Program (grant No. 2023TSYCCX0099) and the National Key R&D Program of China (grant Nos. 2024YFA1611500 and 2022SKA0120102). This work was also partly supported by the Urumqi Nanshan Astronomy and Deep Space Exploration Observation and Research Station of Xinjiang (XJYWZ2303) and the Central Guidance for Local Science and Technology Development Fund (grant No. ZYYD2026JD01). The Nanshan 26-m Radio Telescope (NSRT) is operated by the Xinjiang Astronomical Observatory (XAO), CAS. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. This research has utilized the RATAN-600 data, provided by the Special Astrophysical Observatory of the Russian Academy of Sciences (SAO RAS).
@@ Appendix A Source sample
Appendix B Radio Light curves of Individual Sources Monitored in SMMMAN Program

### v2

@@ I Introduction
A few of the major radio band surveys are briefly introduced below, with emphasis on the importance of multi-frequency radio monitoring projects. The University of Michigan Radio Astronomy Observatory’s (UMRAO; Aller et al. 1985) 26-m paraboloid telescope was devoted to monitoring the flux density and polarization of blazar jets from the mid-1960s until June 2012 (Aller et al., 2017). This facility provided centimeter-band variability over decades at 14.5, 8.0, and 4.8 GHz. The UMRAO data have been used in numerous extensive studies concerning flaring blazars and to acquire information regarding various jet and shock parameters that are not directly observable (Aller et al., 2014; Hughes et al., 2015; Aller et al., 2016) and also in several multiband temporal and spectral studies (e.g. Krishna Mohana et al., 2024; Mohana A et al., 2025). UMRAO observations were also used for the calibration and interpretation of the less frequently sampled Very Long Baseline Array (VLBA) imaging data (Aller et al. 2017, and references therein). It is to be noted that the long-term monitoring data set could help to search for a supermassive black hole binary (SMBHB)
@@
The observations were carried out with the six frequency bands at 1.25, 2.25, 4.7, 8.2, 11.2, and 22.3 GHz, simultaneously (within 1–2 min) when a source moves along the focal line where the receivers are located. These data formed the basis for the first version of the BLcat (Mingaliev et al., 2014), and the latest updated version of this catalog is “RATAN-600 multi-frequency catalog of blazars” (Sotnikova et al., 2022), which includes about 1830 source. Using these catalogs with information of flux densities, spectral indices, variability, and radio luminosity for the large blazar list, several studies to understand the flare activity and synchrotron radio spectra of AGNs were reported (e.g. Mufakharov et al., 2015; Sotnikova et al., 2019; Larionov et al., 2020; Sotnikova et al., 2021; Krishna Mohana et al., 2024; Kunert-Bajraszewska et al., 2025). The Fermi-GST AGN Multi-frequency Monitoring Alliance (F-GAMMA) programme (Fuhrmann et al., 2007; Angelakis et al., 2010; Fuhrmann et al., 2014), conducted at the Effelsberg telescope and IRAM 30 m telescope in Germany, performed nearly monthly broadband radio monitoring of selected blazars from January 2007 to January 2015 (Angelakis et al., 2019a). The monitoring was carried out at multi-frequency from 2.64 to 43 GHz, which includes the frequencies of 2.64, 4.85, 8.35, 10.45, 14.6, 23.05, 32, 43, 86.2, 142.3, and 228.9 GHz. Using initial 3.5 yr data, a cross-correlation analysis was carried out by Fuhrmann et al. (2014) between the radio and -ray light curves of 54 Fermi-bright blazars, with the aim of determining the exact location of the -ray emission region. Fuhrmann et al. (2016) reported a multi-frequency study of AGNs in Fermi era to examine whether the -ray loudness is related to the radio variability. The 2007-2015 multi-frequency radio data published in Angelakis et al. (2019a) is used in several temporal and spectral studies of blazars (e.g. Paliya et al., 2021; Kim et al., 2022; Krishna Mohana et al., 2024).
@@
The Radio Observations of Blazars with INaf telescopes (ROBIN) program has continued to thrive over the years and remains active at frequencies of 8 and 24 GHz, having reached 20 years of operation (Marchili et al., 2025a). The data accumulated through the program have been used in numerous studies and publications, primarily in relation to Whole Earth Blazar Telescope (WEBT) multi-frequency campaigns (e.g. Raiteri et al., 2007; Abdo et al., 2010b; Carnerero et al., 2015; Raiteri et al., 2017, 2024).
The monitoring program has facilitated the examination of how the variability characteristics and SED of blazars evolve over time (Marchili et al. 2025a, and references therein).
At the Aalto University Metsähovi Radio Observatory (MRO, Finland), several hundred AGNs and a few galactic microquasars have been monitored at 22, 37, and 87 GHz since 1980 using the 13.7 m radio telescope (Teraesranta et al., 1998). At present, there are 271 sources in the MRO monitoring program with data at 22 and 37 GHz, where the latter includes observations spanning up to 42 years. Data from this ongoing program have contributed to research on the long-term behavior, statistical properties, source populations, and radio variability models of AGNs. This monitoring has enabled studies of the radio to -ray connection in blazars, as well as radio versus optical behavior and multi-frequency variability models (e.g. Hovatta et al., 2007; Kankkunen et al., 2025). The Submillimeter Array (SMA; Gurwell et al., 2007) in Hawaii, United States, has been monitoring the quasars to determine the flux densities and variability at 230 and 345 GHz since the beginning of 2003 with 437 sources at present. The observations from this program were used in different long-term timing and spectral studies of AGNs (e.g. Villata et al., 2009; Fuhrmann et al., 2014; Jorstad and Marscher, 2016).
@@ II.1 Facility
The SMMAN program is carried out with the NanShan 26-m Radio Telescope (NSRT), which is located on the Nanshan Mountain, with the geographic coordinates of 43°28’17”N, 87°10’41”E, and an altitude of 2080 meters, km southwest of Urumqi city, China. Apart from single-dish scientific observations, NSRT has been involved in a large number of Very Long Baseline Interferometry (VLBI) observations in the last years, as a formal member of the European VLBI Network (EVN). The NSRT has an on-axis, single-beam optical configuration and four receivers operating at L-band (1-2 GHz), S/X-band (2.3/8.6 GHz), C-band (4-8 GHz), and K-band (22-24 GHz). The receivers used in the SMMAN program operate in the C and K bands, centered at 4800 MHz and 23600 MHz, with terminal recording bandwidths of 500 MHz for both. The main technical specifications of the receivers used in the SMMAN program are listed in Table 1.
@@ II.2 Source selection and monitoring schedule
Our goal is to monitor those bright variable blazars, the majority of which have been captured by the Fermi-LAT. Therefore, the initial source sample was established by filtering the Fermi-LAT 3FGL catalog (Acero et al., 2015) and cross-matching with the NRAO VLA Sky Survey (NVSS)222https://www.cv.nrao.edu/nvss/ by putting a limit on flux density at 1.4 GHz (, point-like structure, and source declination (J2000) .
@@
The redshift values were obtained from the NASA/IPAC Extragalactic Database (NED) and the SIMBAD astronomical database (Wenger et al., 2000). Redshift measurements are available for 124 of the 131 sources in our monitored sample, of which 121 are spectroscopic and three are photometric.
@@ II.3 Data Acquisition and Reduction
The data reduction and calibration process includes Gaussian fitting, pointing correction, gain correction, and absolute flux density calibration. Gaussian fitting is used to subtract the baseline of the system temperature drifting and extract the Gaussian-like signals from the sub-scans. Pointing correction is to calibrate the antenna pointing error with each other direction’s offset in the scans. After that, we subsequently conducted atmospheric opacity and antenna gain calibration to further correct the elevation gain effect. Note that the systematic error possibly arising from solar elongation dependence was also accounted for during the pointing correction step. Finally, the absolute flux density was tuned using the calibration factors generated by three primary calibrators, 3C 48, NGC 7027, and 3C 286 (Baars et al., 1977; Ott et al., 1994; Zijlstra et al., 2008; Angelakis et al., 2015). The observed average flux density for these calibrator sources during the SMMAN program is presented in Table 2.
@@
We used the monthly binned -ray flux and spectral index data from the Fermi-LAT Light Curve Repository444https://fermi.gsfc.nasa.gov/ssc/data/access/lat/LightCurveRepository/ (LCR; Abdollahi et al. 2023) in the energy range of GeV, covering the same time period as the radio observations (2016 November 1 to 2024 November 30) presented in this study. For a detailed description of the unbinned maximum-likelihood data analysis, we refer to Abdollahi et al. (2020). In this work, we used the light curve obtained by allowing the spectral slope to vary freely during the fitting process and the fit results of only the convergent likelihood analysis.
The significance of a source detection is evaluated using the test statistic , where is the likelihood of the null model (no source) and is the likelihood of the model that includes the source. In the monthly binned light curve, we only consider those bins with detection significance of (), and the upper limit values (for non-detection bins) were not included. It is important to note that performing a full Fermi-LAT data analysis for 131 SMMAN sources is computationally demanding and time-consuming. Therefore, we used the available light curves from the LCR. These light curves of various -ray sources from this repository have been used in previous research studies (e.g. Di Gesu et al., 2023; Wang et al., 2023).
@@ III Results and Discussion
To depict the importance of the SMMAN monitoring program, a long-term multiband light curve of the BL Lac object S4 0954+65 using the data from the SMMAN, RATAN-600555https://www.sao.ru/blcat/(Mingaliev et al., 2014; Sotnikova et al., 2022), F-GAMMA (Angelakis et al., 2019a, b) and ROBIN (Marchili et al., 2025a, b) monitoring program is shown in Figure 5. The plot clearly shows that the SMMAN program has covered a greater number of observations after 2016 November compared to the other surveys (publicly available), which is important due to the fact that adequate continuous data are essential for long-term temporal and spectral studies, and the detection of QPOs. In addition, a part of the data from the SMMAN program has been used in systematic studies of a few individual blazars in earlier research works (e.g. Kun et al., 2022; Krishna Mohana et al., 2024; Kun et al., 2024; Mohana A et al., 2025), which also showcase the significance of the SMMAN program along with other multiwavelength observations.
@@ III.2 Characteristics of Radio Variability, Spectral index and Luminosity: Relation to AGNs Class
Table 5 of the Appendix A. It should be noted that, among the redshifts used in our calculations, three out of the 124 measurements () are photometric, while the remaining measurements are spectroscopic.
@@ III.3 Gamma-ray and Radio Properties of the SMMAN Sample
Figure 14 shows the radio luminosity at 23.6 GHz as a function of redshift (top panel) and the -ray luminosity as a function of redshift (bottom panel) for the sample of SMMAN sources having radio and -ray luminosities. Clearly, a selection effect is evident: sources with higher luminosities are preferentially detected at higher redshifts in both the radio and -ray bands. FSRQs exhibit higher radio and -ray luminosities and are predominantly found at higher redshifts, whereas BL Lacs tend to have lower radio and -ray luminosities and are generally located at lower redshifts. Radio galaxies span a broader range, extending from the lowest luminosities and redshifts to intermediate values between the FSRQ and BL Lac regimes. Such trend is also noticed for the radio luminosity at 4.8 GHz as a function of redshift.
@@
A linear fit in log-log space gives (Pearson correlation coefficient = 0.68) and (Pearson correlation coefficient = 0.86) for radio luminosities of 4.8 and 23.6 GHz, respectively. Such a correlation could be induced by observational effects, which depend on the source distance (e.g., Böck et al., 2016). This is further supported by Figure 14, which indicates that the observed correlation in Figure 16 is likely largely driven by the fact that our sample is both radio flux-limited and -ray flux-limited. Consequently, a strong correlation is expected as a result of Malmquist bias.
@@ IV Summary and Conclusions
The users using the SMMAN radio data in their publications are kindly requested to acknowledge the source of the data by citing this paper. To use the data and for details, please contact Lang Cui (cuilang@xao.ac.cn).
@@ Acknowledgements
We are thankful to Sándor Frey, Timur Mufakharov, and Junhui Fan for carefully reading the manuscript and providing insightful comments. We also thank the anonymous reviewer for valuable comments, which helped us to improve the manuscript. This work was supported by the Tianshan Talent Training Program (grant No. 2023TSYCCX0099) and the National Key R&D Program of China (grant Nos. 2024YFA1611500 and 2022SKA0120102). This work was also partly supported by the Urumqi Nanshan Astronomy and Deep Space Exploration Observation and Research Station of Xinjiang (XJYWZ2303) and the Central Guidance for Local Science and Technology Development Fund (grant No. ZYYD2026JD01). The Nanshan 26-m Radio Telescope (NSRT) is operated by the Xinjiang Astronomical Observatory (XAO), CAS. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. This research has used the RATAN-600 data, provided by the Special Astrophysical Observatory of the Russian Academy of Sciences (SAO RAS).
@@ Appendix A Source sample
; the number of observations in 4.8 and 23.6 GHz for each source (Col. 7 and 9), and the average flux density calculated over these observations (Col. 8 and 10).
@@
Appendix B Radio Light curves of Individual Sources Monitored in SMMAN Program